OpenStudy (anonymous):

1 + cos(12x) = _____. A. 2sin2(6x) B. 2sin2(12x) C. 2cos2(12x) D. 2cos2(6x)

OpenStudy (anonymous):

\[1+\cos 2\theta=2\cos ^2\theta \]

OpenStudy (midhun.madhu1987):